OpenStudy (anonymous):

1/2 - (-5/9) = A. 6/9 B. 6/18 C. 19/18

OpenStudy (anonymous):

19/18

OpenStudy (anonymous):

\[\Large \begin{array}{l} \frac{1}{2} + \frac{5}{9}\\\\ {\text{LCM}}{\text{ of 2 and 9 = 18, so make both denominators 18:}}\\\\ = \frac{{1(9)}}{{2(9)}} + \frac{{5(2)}}{{9(2)}}\\ = \frac{9}{{18}} + \frac{{10}}{{18}}\\ = \frac{{9 + 10}}{{18}}\\ = \frac{{19}}{{18}} \end{array}\]